A simple pendulum completes one vibration in two seconds. Calculate its length,
-2 when g = 10.0 m s .
1
Expert's answer
2020-08-19T12:56:57-0400

"T=2\pi\sqrt{\frac{l}{g}}\to l=\frac{gT^2}{4\pi^2}=\frac{9.81\cdot2^2}{4\cdot3.14^2}\approx1(m)"